OpenStudy (anonymous):

For a planet moving in an elliptical orbit, why is (1/2)*I*w^2 not the correct expression for its kinetic energy?

OpenStudy (anonymous):

The energy of rotation is given by 1/2Iw², but the total kinetic energy is the energy of rotation plus the energy of rotation. So, the kinetic energy is 1/2m(v²+r²w²), where w is the angular velocity and v is the velocity to increase or decrese distance to the center bdy. If we're on a circular orbit, v²=0.

OpenStudy (mos1635):

 V=ωxR (vectors) ω=RxV/|R^2| (vectors) so ω=R*V*sinθ/R^2 ω=V*sinθ/R